NEW MEASUREMENTS OF THE PROTON MIGRATION TUNNELING SPLITTING IN THE ROTATIONAL SPECTRUM OF C_{2}H_{3}^{+}

\underline{M. CORDONNIER}, Laboratoire de Spectroscopie Hertzienne, U.F.R de Physique, Université des Sciences et Technologies de Lille, Bât. P5, Villeneuve d’Ascq, 59655, France; L. H. COUDERT, Laboratoire de Physique Moléculaire et Applications, C.N.R.S., Université Pierre et Marie Curie, Tour 13, 4, Place Jussieu, 75252 Paris, Cedex 05, France.

New measurements have been carried out in the rotational spectrum of C_{2}H_{3}^{+} using a negative glow discharge. An improved resolution made it possible to determine the value of the tunneling splitting for one more line than in previous studies^{a}^{b}. Microwave data have been analyzed with a formalism based on the Internal-Axis-Method-like approach set-up to treat the water dimer^{c}. The three experimentally resolved splittings were predicted within the experimental resolution around the measured values. Calculations also provided us with more information about the tunneling path along which the three protons migrate^{d}.

aM. Bogey, M. Cordonnier, C. Demuynck and J. L. Destombes, Astrophys. J. \underline{399}, L103 (1992).

bM. Bogey, H. Bolvin, M. Cordonnier, C. Demuynck, J. L. Destombes, R. Escribano and P. C. Gomez, Can. J. Phys. \underline{72}, 967 (1994).

cJ. T. Hougen, J. Mol. Spectrosc. \underline{114}, 395 (1985).

dM. Cordonnier and L. H. Coudert, J. Mol. Spectrosc. submitted (1996).